Total estimated cost of trip $20,000

Assuming that your estimated total cost will grow by 2.5% per year (due to inflation), demonstrate how you would compute the expected future cost of your dream vacation

Suppose that you can invest money every month into a fee-free mutual fund and that this fund is expected to have a 10% nominal annual rate of return. Using your estimated future cost (including inflation) as future value, determine the amount of money you must save each month for the next 10 years (i.e., 120 months) to achieve your goal. Then, determine the monthly amount you must save if you delay your trip for an additional 5 years (that is, you will take the trip 15 years from today = 180 months) instead of 10 years from today. (Note: Be sure to add the 5 additional years of inflation to the estimated future cost.) Write an explanation for your calculations so the reader is completely clear on how you derived your required monthly deposits.

trying to use the excel method... having issues with what to plug in
Answer #1
Estimated vacation cost now $      20,000
Inflation =2.5% per year
Vacation cost after 10 years =20000*(1+2.5%)^10= $25,601.70
So after 10 years I need $25601.70 from the savings of
anniuty.
Formula for future value of Ordinary Annuity :
FV= A [ {(1+k)n-1}/k]
FV = Future annuity value =$25601.7
A = periodical investment
K=interest rate =10% per annum =0.833% per month
N=periods =120 months
So 25601.70= A*[(1+0.000833)^120-1]/0.00833
25601.7=204.798*A
A=$125
So I need to save $125 every month for next 10 years for the vacation.
If I plan the vacation after 15 years , then the vacation cost becomes=
=20000*(1+2.5%)^15=$28965.96
So I need to save $28965.96 in 15 years
Assume monthly saving is A
Formula for future value of Ordinary Annuity :
FV= A [ {(1+k)n-1}/k]
FV = Future annuity value =$28965.96
A = periodical investment
K=interest rate =10% per annum =0.833% per month
N=periods =180 months
A*(1+0.00833)^180-1]/0.00833=28965.96
A*414.318=28965.96
A= $69.91
So I need to save $69.91 per month for next 15 years for the vacation.